April 2013 Reports

To the main

29th April 2013 (Monday)

  • Restriping campaing has started on Saturday
  • T0:
  • T1: RAL (UK sites) Some MC jobs failing to upload from UK sites to different destinations

25th April 2013 (Thursday)

  • Mainly user jobs with some MC ongoing.
  • T0:
  • T1:

22th April 2013 (Tuesday)

  • Mainly user jobs with some MC ongoing.
  • T0:
  • T1:

15th April 2013 (Monday)

  • Apologies for not attending due to LHCb Analysis and Software week
  • Still mainly user jobs
  • T0:
    • Still pb with SLS LFC sensor frown
  • Activities:
    • Re-stripping postponed until beginning of May, will start with latest reprocessed data to take advantage of files still on disk. Then jobs will trigger staging from tape. Tier1s should be prepared for high staging rate then for achieving the rest ripping within 60 days:
Total TB Rate (MB/s)
CERN-RDST 259.4 50
CNAF-RDST 794.6 153
GRIDKA-RDST 643.5 124
IN2P3-RDST 696.8 134
PIC-RDST 201.4 39
RAL-RDST 575.8 111
SARA-RDST 537.7 104

11th April 2013 (Thursday)

  • Mainly user jobs with some MC ongoing.
  • T0:
    • SLS sensor for LHCb LFC flickering every 5 minutes. (RQF0190901)
  • T1:

8th April 2013 (Monday)

  • Mainly user jobs with some MC ongoing.
  • T0:
    • SLS sensor for LHCb LFC flickering every 5 minutes. (RQF0190901)
  • T1:

4th April 2013 (Thursday)

  • Mainly user jobs with some MC ongoing.
  • T0:
    • Nothing to report
  • T1:
    • RAL : Overnight many jobs failed setting up job environment.

2nd April 2013 (Tuesday)

  • Mainly user jobs with some MC ongoing.
  • T0:
    • No SAM tests displayed on the SUM dashboard - solved now (GGUS:92924). Solution not very clear though.
  • T1:
    • RAL : Continuing to have occasional problems with setting up job environment.

28th March 2013 (Thursday)

  • Mainly user jobs with some MC ongoing.
  • T0:
  • T1:
    • PIC: Failed pilots after DT. The jobs submitted had requested 700k HepSpec06.seconds, for some reason they ended up in the short queue (17k sec) (GGUS:92914), no more failures as of yesterday evening, ticket closed
Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2013-09-02 - JoelClosier
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback